Red, Inc., Yellow Corp., and Blue Company each will pay a dividend of $4.89 next year. The growth rate in dividends for all three companies is 5.09 percent. The required return for each company’s stock is 8 percent, 11 percent, and 14 percent, respectively. What is the stock price of the company with the highest stock price
Answer #1

Given for Red Inc.

D1 = $4.89

g = 5.09

Ke = 8%

using Dividend growth model to calculate stock price,

So, P0 = D1*(1+g)/(Ke-g)

So, stock price of Red Inc. = 4.89*1.0509/(0.08-0.0509) = $176.59

Given for Yellow Inc.

D1 = $4.89

g = 5.09

Ke = 11%

using Dividend growth model to calculate stock price,

So, P0 = D1*(1+g)/(Ke-g)

So, stock price of Red Inc. = 4.89*1.0509/(0.11-0.0509) = $86.95

Given for Blue Inc.

D1 = $4.89

g = 5.09

Ke = 14%

using Dividend growth model to calculate stock price,

So, P0 = D1*(1+g)/(Ke-g)

So, stock price of Red Inc. = 4.89*1.0509/(0.14-0.0509) = $57.68

So, the stock price of the company with the highest stock price is for company Red Inc. = $176.59
Answer #2

When given D1, the formula to find current stock value is simply P0 = D1 / (R - g).


P0 = Current Stock Price

D1 = Dividend paid next year = $4.89

R = Required Rate of Return = 0.08 (Red Inc.),

                                                  0.11 (Yellow Corp.),

                                                  0.14 (Blue Company)

g = Growth Rate = 0.0509


So for Red Inc.: P0 = $4.89 / (0.08 - 0.0509) = $168.04

For Yellow Corp.: P0 = $4.89 / (0.11 - 0.0509) = $82.74

For Blue Company: P0 = $4.89 / (0.14 - 0.0509) = $54.88


The company with the highest current stock value should be Red Inc. @ $168.04.